5. A firm produces 100 units of good A at a total cost of $1,500 and separately 200 units of good B at a cost of $2,000. By combining the production of A and B, it is possible to produce the same quantities of A and B respectively at a combined total cost of $2,800. Compute the economies of scope experienced by this firm.

6. The production manager of a clothing manufacturer estimates that the total annual cost of producing men’s suits is given by the equation: C = 5,000 + 4,100Q – 8Q2 + .004Q3. If the market price of suits is constant, what is the shutdown level of output in the short run? What is the minimum price the firm can accept?
Answer #1

Answer)

CA = Cost of producing 100 units good A = $1500

CB = Cost of producing 200 units of good B = $2000

CA+B = Cost of producing A and B together = $2800

Formula for calculation of Economies of Scope is as follows:

S = (CA + CB - CA+B)/(CA+B)

= (1500 + 2000 - 2800)/(2800)

= (3500 - 2800)/(2800)

S  = (700/2800) = 1/4 = 0.25

The cost of producing goods together is 25% less as compared to the cost when the goods are produced seperately.